Kyutiedask10:

The surface area of an oil spill gets 131% larger every day, represented by the function s(x) = (1.31)x − 1. On the first day, it covered an area of 21 square meters. Which function would be used to find the area of the oil spill on the 47th day? A. s(x) = (1.31)21 − 1 B. s(x) = 21(1.31)x − 1 C. s(x) = (1.31)x − 21 D. s(x) = 1.31(21)x − 1

Falconmaster:

I KNOW YOU

Kyutiedask10:

do you?

xxhero782xx:

S(x) = 21(1.31) ^(x -1) Day 1 s(1) = 21 (1.31)^(0) = 21 Day 47 s(47) = 21(1.31)^46 Answer: 21(1.31)^4

Falconmaster:

http://prntscr.com/f1xloj

Kyutiedask10:

That's not one of the answers

Falconmaster:

look at that link i know you

Kyutiedask10:

haha hey

xxhero782xx:

weird